# Conjugacy properties of multivariate unified skew–elliptical distributions

### v2

@@ 3.2 Conjugacy properties of sue distributions in multivariate binary models
thereby recovering the sun prior and skew–normal noise vector considered in Anceschi et al. [4]. Moreover, by Proposition 2 and Remark 2, we have $(\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}\mid{\bm{\beta}})\sim\mbox{SUN}_{n,n}({\bm{D}}_{\bf{y}}{\bf{X}}{\bm{\beta}},\sigma^{2}{\bf{I}}_{n},{\bm{D}}_{\bf{y}}\bar{\alpha}{\bf{I}}_{n},\bm{0},{\bf{I}}_{n})$ which implies
$\mathbb{P}({\bf{y}}\mid{\bm{\beta}})\propto\displaystyle\Phi_{2n}\left(\begin{bmatrix}{\bm{D}}_{\bf{y}}{\bf{X}}{\bm{\beta}}\\ \bm{0}\end{bmatrix};\begin{bmatrix}\sigma^{2}{\bf{I}}_{n}&{\bm{D}}_{\bf{y}}\bar{\alpha}\sigma{\bf{I}}_{n}\\ {\bm{D}}_{\bf{y}}\bar{\alpha}\sigma{\bf{I}}_{n}&{\bf{I}}_{n}\end{bmatrix}\right),\quad\mbox{for all}\ {\bf{y}}\in\{0,1\}^{n},$
which coincides again with the likelihood in Anceschi et al. [4]. As discussed above, such a formulation includes several models of direct interest in practice. For instance, setting $\alpha=0$ yields classical probit regression, whereas replacing $\sigma^{2}{\bf{I}}_{n}$ with a full covariance matrix allows to recover multivariate probit and, for a suitable specification of ${\bf{X}}$, multinomial probit, under both skewed and non–skewed link functions.
As discussed in Example 4 below, these classes of models can be further extended to specific formulations relying on Student’s $t$ and skew–$t$ link functions while preserving conjugacy. As such, the practical impact of Proposition 2 goes beyond the broad class of models studied in Anceschi et al. [4].
Example 4 (Multivariate unified skew–$t$ (sut) conjugacy).
Corollary 2 specializes the conjugacy properties derived in Proposition 2 to the specific context of the sut sub–family presented in Section 2.2.2.
@@
$\mathbb{P}({\bf{y}}\mid{\bm{\beta}})=\frac{\displaystyle T_{n+q}\left(\alpha_{\bm{\beta}}^{-1/2}\begin{bmatrix}{\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}}\\ \bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix};\begin{bmatrix}\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}\\ {\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}^{\top}\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix},\nu+p\right)}{\displaystyle T_{q}(\alpha_{\bm{\beta}}^{-1/2}\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}};\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}},\nu+p)},\quad\mbox{for all}\ {\bf{y}}\in\{0,1\}^{n},$
with $\alpha_{\bm{\beta}}=[\nu+({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})^{\top}\bm{\Omega}_{\bm{\beta}}^{-1}({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})]/(\nu+p)$ and ${\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}}$, $\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}}$, $\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}}$, $\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}$, ${\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}$, $\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}}$ defined as in Proposition 2. Similarly, the resulting posterior distribution is $({\bm{\beta}}\mid{\bf{y}})\sim{{\mbox{SUT}}}_{p,n+q}({\bm{\xi}}_{{\bm{\beta}}\mid{\bf{y}}},\bm{\Omega}_{{\bm{\beta}}\mid{\bf{y}}},{\bm{\Delta}}_{{\bm{\beta}}\mid{\bf{y}}},\bm{\tau}_{{\bm{\beta}}\mid{\bf{y}}},{\bar{\bm{\Gamma}}}_{{\bm{\beta}}\mid{\bf{y}}}{,}\nu)$, with parameters as in Proposition 2.
@@
${\bm{\beta}}\sim\mathcal{T}_{p}({\bm{\xi}}_{\bm{\beta}},\bm{\Omega}_{\bm{\beta}},\nu),\qquad{\bm{\varepsilon}}\sim{{\mbox{SUT}}}_{n,q}(\bm{0},\bm{\Omega}_{{\bm{\varepsilon}}},{\bm{\Delta}}_{{\bm{\varepsilon}}},\bm{0},{\bar{\bm{\Gamma}}}_{{\bm{\varepsilon}}},\nu),$
and hence, under (23), the resulting model for ${\bf{y}}$ coincides with a multivariate binary regression having unified skew–$t$ link function, and Student’s $t$ prior for ${\bm{\beta}}$ uncorrelated with the underlying noise vector ${\bm{\varepsilon}}$. As a direct consequence of Corollary 2, such a formulation yields the likelihood
$\mathbb{P}({\bf{y}}\mid{\bm{\beta}})\propto\displaystyle T_{n+q}\left(\begin{bmatrix}\alpha_{\bm{\beta}}^{-1/2}{\bm{D}}_{\bf{y}}{\bf{X}}{\bm{\beta}}\\ \bm{0}\end{bmatrix};\begin{bmatrix}{\bm{D}}_{\bf{y}}\bm{\Omega}_{\bm{\varepsilon}}{\bm{D}}_{\bf{y}}&\bm{\omega}_{\bm{\varepsilon}}{\bm{D}}_{\bf{y}}{\bm{\Delta}}_{\bm{\varepsilon}}\\ {\bm{\Delta}}^{\top}_{\bm{\varepsilon}}{\bm{D}}_{\bf{y}}\bm{\omega}_{\bm{\varepsilon}}&\bar{\bm{\Gamma}}\end{bmatrix},\nu+p\right),\quad\mbox{for all}\ {\bf{y}}\in\{0,1\}^{n},$
which provides a natural extension to more general settings of classical binary regression with $t$ link function; recall the discussion in Example 2 on the impact of $\alpha_{\bm{\beta}}^{-1/2}$, relative to the standard linear dependence through ${\bf{X}}{\bm{\beta}}$. By setting ${\bm{\Delta}}_{{\bm{\varepsilon}}}=\bm{0}$ within the above formulation, and recalling again Example 2, leads to
@@
which yields a closed–form sut posterior, while further clarifying the direct link with models implemented in practice.
Section 3.3 concludes our analysis by studying sue conjugacy in models for random vectors comprising both fully–observed and dichotomized data. Such a class combines results in Sections 3.13.2 to explore a general set of formulations that extends classical tobit models in both multivariate and skew–elliptical contexts.
@@ 3.3 Conjugacy properties of sue distributions in multivariate censored models
The classes of models studied in Sections 3.13.2 are designed for data that are either all continuous or all discretized. However, in practice, it is also possible to observe vectors comprising a combination of these two types of data. This is the case, for example, when a given continuous variable is fully observed only if its value exceeds a certain threshold. Such a form of censoring is common in several applications and is typically addressed via tobit models and related extensions [3, 29]. Although common implementations rely on Gaussian noise vectors, such a class can be naturally extended to the broader unified skew–elliptical family via the following formulation
${\bf{y}}=[\bar{y}_{1}1(\bar{y}_{1}>0),\ldots,\bar{y}_{n}1(\bar{y}_{n}>0)]^{\top},\quad\mbox{with}\quad\bar{{\bf{y}}}={\bf{X}}{\bm{\beta}}+{\bm{\varepsilon}},$ (25)
@@
Let ${\bf{y}}=({\bf{y}}^{\top}_{1},{\bf{y}}^{\top}_{0})^{\top}$ denote a generic realization from model (25), where ${\bf{y}}_{1}\in\mathbb{R}_{+}^{n_{1}}$ corresponds to the vector of fully–observed data and ${\bf{y}}_{0}=\bm{0}$ comprises the $n_{0}$ censored ones, with $n_{1}+n_{0}=n$. Moreover, assume again $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)})$ and consider the following partition of the parameters
@@
where ${\bm{\varepsilon}}_{1}$ and ${\bm{\varepsilon}}_{0}$ comprise the noise terms associated with the two vectors ${\bf{y}}_{1}$ and ${\bf{y}}_{0}$ in which the generic realization ${\bf{y}}$ is partitioned. Then, when $\bar{{\bf{y}}}$ is defined as in (25), we have that $({\bm{\beta}}^{\top},\bar{\bf{y}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n_{1}+n_{0},q}({\bm{\xi}}^{\dagger},\bm{\Omega}^{\dagger},{\bm{\Delta}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n_{1}+n_{0}+q)}),$ with
@@
where ${\bf{X}}_{1}\in\mathbb{R}^{n_{1}\times p}$ and ${\bf{X}}_{0}\in\mathbb{R}^{n_{0}\times p}$ are the design matrices associated with the two sub–vectors $\bar{{\bf{y}}}_{1}$ and $\bar{{\bf{y}}}_{0}$ of $\bar{{\bf{y}}}=~{}(\bar{{\bf{y}}}^{\top}_{1},\bar{{\bf{y}}}^{\top}_{0})^{\top}$ in (25) — which in turn correspond to the partition ${\bf{y}}=({\bf{y}}^{\top}_{1},{\bf{y}}^{\top}_{0})^{\top}$. Moreover
  1. 1.
@@
    $\begin{split}&{\bm{\xi}}_{{\bm{\beta}}\mid{\bf{y}}}={\bm{\xi}}_{\bm{\beta}}+\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}}),\qquad\bm{\Omega}_{{\bm{\beta}}\mid{\bf{y}}}=\bm{\Omega}_{\beta}-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bm{\beta}}},\qquad\ \bm{\omega}_{{\bm{\beta}}\mid{\bf{y}}}={\mbox{diag}}(\bm{\Omega}_{{\bm{\beta}}\mid{\bf{y}}})^{1/2},\\ &{\bm{\Delta}}_{{\bm{\beta}}\mid{\bf{y}}}=\bm{\omega}_{{\bm{\beta}}\mid{\bf{y}}}^{-1}\left[-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{0}}+\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bf{y}}_{0}}\ \ \ \ \bm{\omega}_{\bm{\beta}}{\bm{\Delta}}_{{\bm{\beta}}}-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\omega}_{{\bf{y}}_{1}}{\bm{\Delta}}_{{\bf{y}}_{1}}\right]\bm{\gamma}_{{\bm{\beta}}\mid{\bf{y}}}^{-1},\quad Q_{{\bf{y}}_{1}}({\bf{y}}_{1})=({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}})^{\top}\bm{\Omega}_{{\bf{y}}_{1}}^{-1}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}}),\\ &\bm{\tau}_{{\bm{\beta}}\mid{\bf{y}}}=\bm{\gamma}_{{\bm{\beta}}\mid{\bf{y}}}^{-1}\begin{bmatrix}-{\bm{\xi}}_{{\bf{y}}_{0}}-\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}}),\\ \bm{\tau}+{\bm{\Delta}}^{\top}_{{\bf{y}}_{1}}\bar{\bm{\Omega}}^{-1}_{{\bf{y}}_{1}}\bm{\omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}})\end{bmatrix},\ \ \bar{\bm{\Gamma}}_{{\bm{\beta}}\mid{\bf{y}}}=\bm{\gamma}_{{\bm{\beta}}\mid{\bf{y}}}^{-1}\begin{bmatrix}\bm{\Omega}_{{\bf{y}}_{0}}-\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bf{y}}_{0}}&-\bm{\omega}_{{\bf{y}}_{0}}{\bm{\Delta}}_{{\bf{y}}_{0}}+\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\omega}_{{\bf{y}}_{1}}{\bm{\Delta}}_{{\bf{y}}_{1}}\\ (-\bm{\omega}_{{\bf{y}}_{0}}{\bm{\Delta}}_{{\bf{y}}_{0}}+\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\omega}_{{\bf{y}}_{1}}{\bm{\Delta}}_{{\bf{y}}_{1}})^{\top}&\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}_{{\bf{y}}_{1}}\bar{\bm{\Omega}}_{{\bf{y}}_{1}}^{-1}{\bm{\Delta}}_{{\bf{y}}_{1}}\end{bmatrix}\bm{\gamma}_{{\bm{\beta}}\mid{\bf{y}}}^{-1},\end{split}$

### v3

@@ 3.2 Conjugacy properties of SUE distributions in multivariate binary models
thereby recovering the SUN prior and skew-normal noise vector considered in Anceschi et al. [4]. Moreover, by Proposition 2 and Remark 2, we have $(\bar{{\bf{y}}}_{{\bm{D}}_{\bf{y}}}\mid{\bm{\beta}})\sim\mbox{SUN}_{n,n}({\bm{D}}_{\bf{y}}{\bf{X}}{\bm{\beta}},\sigma^{2}{\bf{I}}_{n},{\bm{D}}_{\bf{y}}\bar{\alpha}{\bf{I}}_{n},\bm{0},{\bf{I}}_{n})$ which implies
$\mathbb{P}({\bf{y}}\mid{\bm{\beta}})\propto\displaystyle\Phi_{2n}\left(\begin{bmatrix}{\bm{D}}_{\bf{y}}{\bf{X}}{\bm{\beta}}\\ \bm{0}\end{bmatrix};\begin{bmatrix}\sigma^{2}{\bf{I}}_{n}&{\bm{D}}_{\bf{y}}\bar{\alpha}\sigma{\bf{I}}_{n}\\ {\bm{D}}_{\bf{y}}\bar{\alpha}\sigma{\bf{I}}_{n}&{\bf{I}}_{n}\end{bmatrix}\right),$
for all ${\bf{y}}\in\{0,1\}^{n}$, which coincides again with the likelihood in Anceschi et al. [4]. As discussed above, such a formulation includes several models of direct interest in practice. For instance, setting $\alpha=0$ yields classical probit regression, whereas replacing $\sigma^{2}{\bf{I}}_{n}$ with a full covariance matrix allows to recover multivariate probit and, for a suitable specification of ${\bf{X}}$, multinomial probit, under both skewed and non-skewed link functions.
As discussed in Example 4 below, these classes of models can be further extended to specific formulations relying on Student’s $t$ and skew-$t$ link functions while preserving conjugacy. As such, the practical impact of Proposition 2 goes beyond the broad class of models studied in Anceschi et al. [4].
Example 4 (Multivariate unified skew-$t$ (SUT) conjugacy).
Corollary 2 specializes the conjugacy properties derived in Proposition 2 to the specific context of the SUT sub-family presented in Section 2.2.2.
@@
$\mathbb{P}({\bf{y}}\mid{\bm{\beta}})=\frac{\displaystyle T_{n+q}\left(\alpha_{\bm{\beta}}^{-1/2}\begin{bmatrix}{\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}}\\ \bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix};\begin{bmatrix}\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}\\ {\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}^{\top}\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}&\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}}\end{bmatrix},\nu+p\right)}{\displaystyle T_{q}(\alpha_{\bm{\beta}}^{-1/2}\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}};\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}},\nu+p)},$
for all ${\bf{y}}\in\{0,1\}^{n},$ with $\alpha_{\bm{\beta}}=[\nu+({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})^{\top}\bm{\Omega}_{\bm{\beta}}^{-1}({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})]/(\nu+p)$, and ${\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}}$, $\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}}$, $\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}}$, $\bm{\omega}_{{\bf{y}}\mid{\bm{\beta}}}$, ${\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}}$, $\bar{\bm{\Gamma}}_{{\bf{y}}\mid{\bm{\beta}}}$ as in Proposition 2. The resulting posterior is $({\bm{\beta}}\mid{\bf{y}})\sim{{\mbox{SUT}}}_{p,n+q}({\bm{\xi}}_{{\bm{\beta}}\mid{\bf{y}}},\bm{\Omega}_{{\bm{\beta}}\mid{\bf{y}}},{\bm{\Delta}}_{{\bm{\beta}}\mid{\bf{y}}},\bm{\tau}_{{\bm{\beta}}\mid{\bf{y}}},{\bar{\bm{\Gamma}}}_{{\bm{\beta}}\mid{\bf{y}}}{,}\nu)$, with parameters as in Proposition 2.
@@
and hence, under (23), the resulting model for ${\bf{y}}$ coincides with a multivariate binary regression having unified skew-$t$ link function, and Student’s $t$ prior for ${\bm{\beta}}$ uncorrelated with the underlying noise vector ${\bm{\varepsilon}}$. As a direct consequence of Corollary 2, such a formulation yields the likelihood
$\mathbb{P}({\bf{y}}\mid{\bm{\beta}})\propto\displaystyle T_{n+q}\left(\begin{bmatrix}\alpha_{\bm{\beta}}^{-1/2}{\bm{D}}_{\bf{y}}{\bf{X}}{\bm{\beta}}\\ \bm{0}\end{bmatrix};\begin{bmatrix}{\bm{D}}_{\bf{y}}\bm{\Omega}_{\bm{\varepsilon}}{\bm{D}}_{\bf{y}}&\bm{\omega}_{\bm{\varepsilon}}{\bm{D}}_{\bf{y}}{\bm{\Delta}}_{\bm{\varepsilon}}\\ {\bm{\Delta}}^{\top}_{\bm{\varepsilon}}{\bm{D}}_{\bf{y}}\bm{\omega}_{\bm{\varepsilon}}&\bar{\bm{\Gamma}}\end{bmatrix},\nu+p\right),$
for all ${\bf{y}}\in\{0,1\}^{n}$, which provides a natural extension to more general settings of classical binary regression with $t$ link function; recall the discussion in Example 2 on the impact of $\alpha_{\bm{\beta}}^{-1/2}$, relative to the standard linear dependence through ${\bf{X}}{\bm{\beta}}$. By setting ${\bm{\Delta}}_{{\bm{\varepsilon}}}=\bm{0}$ within the above formulation, and recalling again Example 2, leads to
@@
which yields a closed-form SUT posterior, while further clarifying the direct link with models implemented in practice.
Section 3.3 concludes our analysis by studying SUE conjugacy in regression models for random vectors comprising both fully-observed and dichotomized data. Such a class combines results in Sections 3.13.2 to explore a general set of formulations that extends classical tobit models in both multivariate and skew-elliptical contexts.
@@ 3.3 Conjugacy properties of SUE distributions in multivariate censored models
The classes of models studied in Sections 3.13.2 are designed for data that are either all continuous or all discretized. However, in practice, it is also possible to observe vectors comprising a combination of these two types of data. This is the case, for example, when a continuous variable is fully observed only if its value exceeds a certain threshold. Such a form of censoring is common in several applications and is typically addressed via tobit models and related extensions [3, 29]. Although common implementations rely on Gaussian noise vectors, such a class can be naturally extended to the broader unified skew-elliptical family via the following formulation
${\bf{y}}=[\bar{y}_{1}1(\bar{y}_{1}>0),\ldots,\bar{y}_{n}1(\bar{y}_{n}>0)]^{\top},\qquad\qquad\bar{{\bf{y}}}={\bf{X}}{\bm{\beta}}+{\bm{\varepsilon}},$ (25)
@@
Let ${\bf{y}}=({\bf{y}}^{\top}_{1},{\bf{y}}^{\top}_{0})^{\top}$ denote a generic realization from model (25), where ${\bf{y}}_{1}\in\mathbb{R}_{+}^{n_{1}}$ corresponds to the vector of fully-observed data and ${\bf{y}}_{0}=\bm{0}$ comprises the $n_{0}$ censored ones, with $n_{1}+n_{0}=n$. Moreover, assume again $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)})$ and consider the following partition of the parameters
@@
where ${\bm{\varepsilon}}_{1}$ and ${\bm{\varepsilon}}_{0}$ comprise the noise terms associated with the two vectors ${\bf{y}}_{1}$ and ${\bf{y}}_{0}$ in which the generic realization ${\bf{y}}$ is partitioned. Then, when $\bar{{\bf{y}}}$ is defined as in (25), we have $({\bm{\beta}}^{\top},\bar{\bf{y}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n_{1}+n_{0},q}({\bm{\xi}}^{\dagger},\bm{\Omega}^{\dagger},{\bm{\Delta}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n_{1}+n_{0}+q)})$ with
@@
where ${\bf{X}}_{1}\in\mathbb{R}^{n_{1}\times p}$ and ${\bf{X}}_{0}\in\mathbb{R}^{n_{0}\times p}$ denote the two design matrices associated with the two sub-vectors $\bar{{\bf{y}}}_{1}$ and $\bar{{\bf{y}}}_{0}$ of $\bar{{\bf{y}}}=~{}(\bar{{\bf{y}}}^{\top}_{1},\bar{{\bf{y}}}^{\top}_{0})^{\top}$ in (25), which in turn correspond to the partition ${\bf{y}}=({\bf{y}}^{\top}_{1},{\bf{y}}^{\top}_{0})^{\top}$. Moreover
  1. (a)
@@
    $\begin{split}&{\bm{\xi}}_{{\bm{\beta}}\mid{\bf{y}}}={\bm{\xi}}_{\bm{\beta}}+\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}}),\qquad\bm{\Omega}_{{\bm{\beta}}\mid{\bf{y}}}=\bm{\Omega}_{\beta}-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bm{\beta}}},\qquad\ \bm{\omega}_{{\bm{\beta}}\mid{\bf{y}}}={\mbox{diag}}(\bm{\Omega}_{{\bm{\beta}}\mid{\bf{y}}})^{1/2},\\ &{\bm{\Delta}}_{{\bm{\beta}}\mid{\bf{y}}}=\bm{\omega}_{{\bm{\beta}}\mid{\bf{y}}}^{-1}\left[-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{0}}+\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bf{y}}_{0}}\ \ \ \ \bm{\omega}_{\bm{\beta}}{\bm{\Delta}}_{{\bm{\beta}}}-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\omega}_{{\bf{y}}_{1}}{\bm{\Delta}}_{{\bf{y}}_{1}}\right]\bm{\gamma}_{{\bm{\beta}}\mid{\bf{y}}}^{-1},\quad Q_{{\bf{y}}_{1}}({\bf{y}}_{1})=({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}})^{\top}\bm{\Omega}_{{\bf{y}}_{1}}^{-1}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}}),\\ &\bm{\tau}_{{\bm{\beta}}\mid{\bf{y}}}=\bm{\gamma}_{{\bm{\beta}}\mid{\bf{y}}}^{-1}[(-{\bm{\xi}}_{{\bf{y}}_{0}}-\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}}))^{\intercal}\ \ \ (\bm{\tau}+{\bm{\Delta}}^{\top}_{{\bf{y}}_{1}}\bar{\bm{\Omega}}^{-1}_{{\bf{y}}_{1}}\bm{\omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}}))^{\intercal}]^{\intercal},\\ &\bar{\bm{\Gamma}}_{{\bm{\beta}}\mid{\bf{y}}}=\bm{\gamma}_{{\bm{\beta}}\mid{\bf{y}}}^{-1}\begin{bmatrix}\bm{\Omega}_{{\bf{y}}_{0}}-\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bf{y}}_{0}}&-\bm{\omega}_{{\bf{y}}_{0}}{\bm{\Delta}}_{{\bf{y}}_{0}}+\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\omega}_{{\bf{y}}_{1}}{\bm{\Delta}}_{{\bf{y}}_{1}}\\ (-\bm{\omega}_{{\bf{y}}_{0}}{\bm{\Delta}}_{{\bf{y}}_{0}}+\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\omega}_{{\bf{y}}_{1}}{\bm{\Delta}}_{{\bf{y}}_{1}})^{\top}&\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}_{{\bf{y}}_{1}}\bar{\bm{\Omega}}_{{\bf{y}}_{1}}^{-1}{\bm{\Delta}}_{{\bf{y}}_{1}}\end{bmatrix}\bm{\gamma}_{{\bm{\beta}}\mid{\bf{y}}}^{-1},\end{split}$